Suppose that Tata's Consultant estimated the following regression equation for Indica Automobiles:


Q1 = 100,000 - 100P1 +2000N +50I + 30 Pm -1000 Pp +3A +40,000 Pi


Q1 = Quantity demanded per year of Indica Automobiles

P1 = price of automobiles, in dollars

N = Population of India, im millions

I = Per Capita Disposable Income

Pm = Price of Maruti Automobiles

Pp = Real Price of petrol, in cents per gallon

A = adevertising expenditures by Indica, in dollars per year

Pi = credit incentives to purchase Indica, in percentage points below the rate of interest on borrowing in the absence of incentives


(a) Indicate the change in the number of Indicas purchased per year (Q1) for each unit change in the independent or explanatory variables


(b) Find the of Q1 if the average value of P1 = $ 9,000, N = 200 million, I = $ 10,000, Pm = $ 8,000, Pp = 80 cents, A = $ 200,000, and it Pi = 1



1
Expert's answer
2021-04-16T07:28:17-0400

Q1 = 100,000 - 100P1 + 2000N + 50I + 30Pm - 1000 Pp + 3A + 40,000 Pi.

(a) The number of Indicas purchased per year (Q1) will increase when N, I, Pm, A or Pi increase, and will decrease when P1 or Pp decrease.

(b) If P1 = $9,000, N = 200 million, I = $10,000, Pm = $8,000, Pp = 80 cents, A = $200,000, and it Pi = 1, then:

Q1 = 100,000 - 100×9,000 + 2,000×200 + 50×10,000 + 30×8,000 - 1000×0.8 + 3×200,000 + 40,000×1 = 900 units.